The number of hours between servicing required for a particular jet ski engine is exponentially distributed with a mean time of 184 hours. Determine the probability that a randomly selected engine:

a) will run for less than 200 hours before servicing is needed. (5 pts)

b) will run more than 240 hours before servicing is needed. (5 pts)

Do you work in Excel. Use the =expon.dist( ) function
Answer #1

Ans:

Given that

mean time=184

P(T<=t)=1-e-t/184

a)

P(T<=200)=1-e-200/184=0.6628

b)

P(T>240)=e-240/184=0.2713